There are many brown spots where the grass died and tired of moving you decided to install artificial grass. The area is 15 ft by 10 ft. The artificial grass is conveniently sold in rolls that are 5 yards wide and will be cut to the length needed (to the nearest foot). The fake grass is $54 per linear foot. (For example, you could buy 4 linear feet of grass that is 5 yards wide for 4 x $54 = $216).
What is the amount of artificial grass that is needed and what will be your cost?
thank you for the help.

Answer by CubeyThePenguin(3113)

area of yard = (15)(10) = 150 ft^2 = 5 5/9 yd^2

cost of grass: (5 yd)($54/ft) = $810/yd^2

amount of money to spend =(5 5/9 yd^2) * $810/yd^2 = $4500